Falcon Networks

Falcon Networks is a leading semiconductor company with operations in 17 different countries. Information about the company's taxes appears below:

Falcon Networks

Components of Income Tax Expense

(in millions)

2012

2011

Current - Federal

$ 55.65

$ 47.52

- Foreign

83.85

78.95

- State and Local

14.69

12.5

Total Current

$154.19

$138.97

Deferred - Federal

$ 30.28

$ 42.90

- Foreign

23.89

14.58

Total Deferred

$ 54.17

$ 57.48

Total Income Tax Expense

$208.36

$196.45

Note: Falcon Networks has no current liability at year-end with respect to total current taxes.

Components of Income before Taxes

2012

2011

United States

$256.35

$253.68

Foreign

236.85

198.85

Total

$493.20

$452.53

1.Based on the information provided by Falcon Networks how much cash did income taxes use during 2012?

Group of answer choices

$154.19 million

$54.17 million

$208.36 million

$284.84 million

2.Using the information provided by Falcon Networks, determine the federal effective tax rate for 2012.

Group of answer choices

33.52%

35.00%

42.25%

45.49%

3.

Using the information provided by Falcon Networks, determine the foreign effective tax rate for 2012.

Group of answer choices

33.52%

35.00%

42.25%

45.49%

4.

Using the information provided by Falcon Networks determine the combined effective tax rate for 2012.

Group of answer choices

33.52%

35.00%

42.25%

45.49%
Answer #1

1. Falcon Networks used $ 154.19 million cash during income taxes use in 2012

2. Federal effective tax rate for 2012. = 55.65+30.28/256.35 = 33.52%

3. Foreign effective tax rate for 2012 = 107.74 / 236.85 = 45.49%

4. combined effective tax rate for 2012. = 208.36 / 493.2 = 42.25%

Hope this helps, if not please let know in comments. We expert here put lots of effort to give correct and quality answer to students. Please mark the answer as helpful for the efforts put, it will mean alot. Thanks